OpenStudy (anonymous):

Write the equation in slope intercept form of the line that has a slope of -3 and contains the point (1,1)

OpenStudy (anonymous):

Slope is y =mx+b. Where b is the y-intercept, and m is the slope. Substitute your slope to be m: y=-3x+b Now, substitute (1,1) in to get your y-intercept. 1 = -3(1) +b 1 = -3 +b To find b, move all to the other side. 1 + 3 = b 4 = b Therefore: y=-3x+4